The sum of an integer and its square is 30. Find the integer. Please help
Answer by SadieKhan(31)

Let x be the integer and x^2 be its square
by requirement
x^2+x=30
solving
x^2+x-30=0
factorizing
X^2+6x-5x-30=0
x(x+6)-5(x+6)=0
x-5=0,x+6=0
this implies
x=5,-6
So it can be either a 5 or a -6.